I have two problems I need help with. WIll you please help me.The problems are as following:
write each equation in slope-intercept form. Then identy the slope and the y intercept for each line. ex. 2x-y=12
and find an equation of the line that passes through the given point and has the given slope. Express your answer in slope-intercept form. (8,0), m=-3

Answer by ankor@dixie-net.com(22740)

write each equation in slope-intercept form.
Then identify the slope and the y intercept for each line. ex. 2x-y=12
Put the equation into the slope/intercept form (y - mx+b)
2x - y = 12
subtract 2x from both sides
-y = -2x + 12
y has to be positive, multiply both sides by -1
y = 2x - 12
the slope = 2, and the y intercept = -12 (occurs when x=0)
:
find an equation of the line that passes through the given point and has the given slope.
Express your answer in slope-intercept form. (8,0), m=-3
:
Use the point slope form: y - y1 = m(x - x1), where:
m = -3; x1 = 8; y1 = 0
y - 0 = -3(x - 8)
y = -3x + 24; the slope intercept form of the line (slope = -3, y intercept = 24)
:
:
You can confirm this by substituting 8 for x and seeing that y = 0